### v1

@@ Abstract
PLA-Serve identifies and disaggregates requests with different prompt lengths in LLM serving to reduce TTFT latency. While recent systems have decoupled the prefill and decode stages to improve throughput, they still rely on unified scheduling policies that fail to adapt to heterogeneous workload characteristics. We observe that prompt-length variations lead to distinct performance bottlenecks, motivating an adaptive scheduling strategy. PLA-Serve disaggregates multi-turn long-prefill requests from short-prefill ones and introduces a length-aware smart batching mechanism for short-prefill workloads. It adopts a dual-queue design that supports temporal disaggregation on a single prefill instance or spatial disaggregation across multiple instances. For short-prefill batches, a batch waiting window and CUDA Graph-based clustering mitigate interference from heterogeneous computation, reducing batching delay and lowering average latency. In real multi-turn workloads, PLA-Serve reduces prefill latency by over 30% compared to vanilla SGLang under prefill–decode disaggregation, and further decreases SLO violations by 28% in multi-instance deployments with vanilla data-parallel configuration. Compared to the SGLang router with load balancing, it further lowers SLO violations by 12% in multi-GPU settings. Under high concurrency and mixed-request scenarios, PLA-Serve improves request throughput by 35% serving Qwen2.5-32B model for prefill instance, demonstrating its effectiveness in optimizing heterogeneous LLM serving workloads.
@@ 1 Introduction
Modern LLM serving stacks (e.g., vLLM kwon2023efficientmemorymanagementlarge, SGLang zheng2024sglangefficientexecutionstructured) combine prefill-decoding (PD) disaggregation zhong2024distservedisaggregatingprefilldecoding with continuous batching to meet low-latency, high-concurrency service-level objectives (SLOs). The Prefill stage (first-token computation) is largely compute-bound, while the decoding stage (auto-regressive generation) is memory-bound. PD disaggregation decouples these two phases across separate instances to avoid cross-phase contention. On the prefill side, concurrent requests are batched to raise GPU utilization. However, we show that this separation and batching are insufficient: even with PD, interference still persists within the prefill stage when long, compute-bound prefills are mixed with short, memory-bound prefills/re-prefills.
Figure 1: P90 TTFT of long-prefill requests under varying concurrency levels for long and short requests. The long-prefill requests have more than 1K tokens, while the short ones have fewer than 64 tokens. We concurrently run them on a single H200 GPU and serve by Qwen2.5-32B qwen2025qwen25technicalreport. The dashed lines indicate the latency when only long-prefill requests are served.
@@
Re‑prefill denotes the repeated prefill in multi‑turn sessions where the model extends an existing context by combining new tokens with cached KV states. It is common in chatbots dam2024completesurveyllmbasedai, tool-using agents wölflein2025llmagentsmakingagent, RAG, speculative decoding leviathan2023fastinferencetransformersspeculative, and token routing she2025tokenlevelroutinginference, and is typically memory-bound (dominated by KV-cache reads/writes rather than large GEMMs). Figure 2 illustrates the token length distribution of a real-world trace, LMsys-Chat-1M zheng2024lmsyschat1mlargescalerealworldllm, which is collected from real human-AI conversations. We observe that most prompts are short (<256 tokens), while long-context requests (>1K tokens) are relatively rare. This indicates that production workloads are dominated by short prefills/re-prefills, which are memory-bound. Mixing them with long, compute-bound prefills in unified batches will also induce compute-memory interference: short prefills/re-prefills wait behind long GEMMs and time-to-first-token (TTFT) spikes, and long prefills lose effective FLOPs due to heavy KV traffic from short jobs. Figure 1 confirms this issue: mixing long and short prefills does significantly increase long-prefill latency, and this contention becomes more severe as concurrency rises.
@@
Although prior systems have recognized the resource contention between compute-bound and memory-bound workloads, they focus on coordinating prefill and decode: (1) decode‑priority schedulerskwon2023efficientmemorymanagementlarge prioritize the decode phase to minimize per-token latency; (2) prefill-first schedulers prioritize the prefill phase and use continuous batching yu2022orca to improve throughput; (3) stall-free chunked prefill agrawal2023sarathiefficientllminference splits long prefill into chunks and interleaves them with decode, so long prefills don’t stall ongoing decodes; and (4) PD disaggregation zhong2024distservedisaggregatingprefilldecoding; jin2024pdserveservingdisaggregatedlarge; hu2024inferenceinterferencedisaggregatellm; strati2024dejavukvcachestreamingfast, which dominates in modern serving systems, runs prefill and decode on separate instances. These advances alleviate cross‑phase contention but implicitly assume all prefills are compute‑bound long sequences, overlooking that short (e.g., <256 tokens), memory‑bound prefills/re‑prefills could dominate real‑world multi-turn serving workloads.
@@
To address these challenges, we propose PLA-Serve, a length-aware LLM serving system that explicitly disaggregates and optimizes long-prefill and short-prefill workloads within the prefill stage. PLA-Serve maintains two separate prefill pools at runtime and performs batch disaggregation to isolate long and short prefill requests, completely eliminating their mutual interference. For short-prefill requests, PLA-Serve introduces a dynamic waiting window in the scheduler and bucketizes requests by input length, making them more uniform and enabling larger batch sizes. This reduces batch-launch overhead and, combined with CUDA Graph-based execution, further accelerates processing and improves throughput. For SLO-serving scenarios, we design an SLO-aware scheduler that balances the trade-off between the waiting window and throughput.
@@
In multi-instance deployments, the scheduler dynamically adjusts each instance’s workload type based on real-time load, achieving load balancing across spatially disaggregated prefill instances. This adaptive strategy resembles resource allocation problems in deep learning systems, like in qiao2021pollux.
@@
In addition, LLM serving systems typically support three modes: mix and PD temporal/spatial disaggregation. Beyond these, our PLA-Serve introduces a fourth mode:
@@
It is worth noting that our PLA-Serve remains compatible with existing PD disaggregation architectures. Overall, our main contributions are summarized as follows:
@@ 2.1 Compute-Memory Boundary for (Re-)Prefills
Empirical profiling across advanced hardware generations (A100, H100, and H200) and LLMs ranging from 7B to 32B parameters shows that this transition typically occurs between 150 and 512 tokens yuan2024llminferenceunveiledsurvey; zhong2024distservedisaggregatingprefilldecoding, depending on model architecture, kernel implementation, and batch configuration. Prefills shorter than this range tend to be memory-bound and limited by KV-cache I/O, while longer ones are dominated by GEMM throughput.
@@ 2.2 Exploring Interference Between Long-Short Prefills/Re-prefills
We model the interference between long-prefill and short-prefill (or re-prefill) requests using a standard M/G/1 (Markovian/General/1-server queue) first-come, first-served (FCFS) queuing model meini1998solving. Unlike prior PD scheduling analyses that focused on prefill-decode contention, we examine the intra-prefill interference caused by mixing compute-bound and memory-bound jobs within continuous batching.
@@
Using the Pollaczek-Khinchine (P-K) result neuts1986generalizations for M/G/1 queues, the mean waiting time is
@@ 2.3 Uncovering the Sources of Long-Short Prefill Mixing
General-purpose LLM services must handle a wide spectrum of tasks. As shown in Table 1, daily chat and creative ideation are typical short-prompt tasks, while speculative decoding and token routing produce high-frequency, short re-prefills chatterji2025people. In contrast, long-document QA and autonomous agent workflows correspond to long-context prefills. In practice, these streams interleave over time, leading to long-short mixing within prefill workloads.
Most existing systems schedule requests in an FCFS fashion, packing them into unified batches. Many deploy multi-queue variants: continuous or rolling batching (e.g., vLLM kwon2023efficientmemorymanagementlarge, TGI huggingface_tgi_docs_2024) treats prefill and decode as distinct phases, applying FCFS-style admission under token/KV limits and optional priorities or aging policies. With chunked prefill (e.g., Sarathi-Serve agrawal2023sarathiefficientllminference), vLLM prioritizes decode and may co-batch prefills with decode. In-flight batching (e.g., TensorRT-LLM tensorrt_llm) runs distinct context and generation engines, each maintaining its own ready queue and often prioritizing generation. Moreover, many LLM gateways expose service tiers that prioritize certain request classes while enforcing token-based budgets using fixed or sliding windows (e.g., OpenAI openai_pricing, Anthropic anthropic_pricing, Envoy envoy_gateway, Kong kong_gateway, APISIX apisix_gateway, and Cloudflare cloudflare_api).
@@
The most related line of work is length bucketing, which groups requests by predicted sequence length into size-homogeneous buckets to reduce padding and improve throughput (e.g., Multi-Bin Batching guldogan2024multi, BucketServe zheng2025bucketserve). However, these methods only optimize intra-batch length variance; they do not disaggregate prefills versus re-prefills nor address the compute-memory interference we identify.
3 Prefill-Length-Aware (PLA) Serve
@@
We develop PLA-Serve to mitigate the interference between long-prefill and short-prefill requests in multi-turn LLM serving. The interference stems from two major factors: (1) their heterogeneous computation characteristics, and (2) head-of-line blocking caused by unified batching. Section 3.1 introduces the strategies we design to optimize high-concurrency short-prefill workloads, while Section 3.2 shows our queue- and instance-level disaggregation mechanism that isolates these request types to reduce interference. PLA-Serve is built upon the prefill instance in the PD disaggregation architecture and extends it with a finer-grained disaggregation design within the prefill stage.
@@ 3.1 Short Prefill Optimization
During auto-regressive serving, the majority of end-to-end latency typically arises from the decode stage. Consequently, most existing optimization efforts (e.g., PD disaggregation zhong2024distservedisaggregatingprefilldecoding, CUDA Graph acceleration harish2007accelerating, and router-based load balancing across decoding instances hu2024routerbenchbenchmarkmultillmrouting; jain2025intelligentrouterllmworkloads; stripelis2024tensoropera; jitkrittum2025universal) have been designed for the decode phase. However, as the diversity of LLM workloads grows (e.g., agent decision-making, chain-of-thought reasoning, and multi-turn task planning), the prefill stage has become an increasingly significant bottleneck, yet its optimization potential remains largely overlooked. Despite this trend, optimization for short and multi-turn prefills remains unexplored.
@@
Graph capture and bucketization. From the characteristics of intensive re-prefill workloads, speculative decoding and token-level routing—although not conventional PD inference—generate numerous short re-prefill requests. Drawing inspiration from EAGLE-2’s speculative decoding optimization li2024eagle2fasterinferencelanguage; li2025eagle3scalinginferenceacceleration, PLA-Serve pre-defines a grid of power-of-two prompt-length-batch-size buckets (e.g., $L\in\{8,16,32,64,128,256\}$ and $B\in\{1,2,4,8,16,32,64\}$) gao2023tbdb. At system initialization, a CUDA Graph is captured for each bucket under the assumption of fixed operator topology and variable memory addresses. During inference, each short-prefill request is padded to the nearest bucket and grouped with others sharing the same $(L,B)$ configuration, thereby maximizing graph reuse with negligible memory overhead.
Graph-aware memory-based batching. To maximize the benefit of CUDA Graphs for short-prefill workloads, PLA-Serve optimizes batching with two goals: (i) reducing Graph launch frequency and (ii) increasing the reuse rate of large-batch Graphs. These are achieved through modestly extended waiting windows and graph fusion. Under high concurrency, slightly delaying batch formation allows more short-prefill requests to accumulate, improving overall efficiency when the saved launch overhead outweighs the waiting cost. Figure 5 shows the latency-throughput trade-off under different window settings.
@@
While current serving systems typically adopt a memory-constrained batching policy (i.e., aggregating requests until total tokens reach the GPU memory limit), PLA-Serve enhances this approach with graph awareness. During short-prefill batching, requests are grouped under the memory budget and aligned to the nearest captured Graph shape.
@@ 3.2 Long-Short Prefill Disaggregation
To fundamentally eliminate the interference between long- and short-prefill requests discussed above, we adopt a design philosophy inspired by PD disaggregation, which further disaggregates long-prefill (LP) and short-prefill (SP) requests. However, unlike PD disaggregation, where the prefill and decode stages exhibit strong temporal dependencies and KV-cache transfers, the two types of tasks are merely mutually exclusive in our LP/SP disaggregation, resulting in fewer constraints in the scheduling objectives. Consequently, practical PD instance scheduling must account for the capacity coordination between the prefill and decode cluster, as well as the effective interconnect bandwidth when designing resource allocation strategies, but our design provides a larger design space for scheduling strategies that can adapt to the physical compute resources and workload characteristics. To this end, PLA-Serve implements two complementary schedulers: a temporal disaggregation scheduler for single-instance prefill execution, and a spatial disaggregation scheduler for multi-instance prefill coordination. Figure 4 presents the system overview.
Disaggregating prefill execution eliminates direct interference between compute-bound long-prefill and memory-bound short-prefill tasks; however, static separation alone cannot accommodate dynamic workload variations. In real-world deployments, the ratio of long to short requests fluctuates over time, and requests within each category exhibit heterogeneous lengths and deadlines. To address this, PLA-Serve introduces a hierarchical scheduling layer: a temporal disaggregation scheduler is employed within each single prefill instance to manage intra-instance prioritization, while a spatial disaggregation scheduler operates across multiple prefill instances to coordinate inter-instance resource allocation.
It is worth noting that the disaggregation design further amplifies the benefits of CUDA Graphs for short-prefill workloads. In mixed long/short prefill instances, a unified queue containing both request types limits the ability of short-prefill requests to form large, graph-aligned batches. In contrast, by maintaining two independent queues under the disaggregated design, PLA-Serve can determine at request arrival whether CUDA Graph execution should be applied, thereby reducing batching delay and minimizing shape heterogeneity. In mixed queues, the large length disparity between long and short requests leads to excessive padding, lowering the Graph shape hit rate and GPU memory efficiency. After LP/SP separation, requests within each instance exhibit a more concentrated length distribution, improving both Graph reuse and throughput. As a result, the system can achieve higher CUDA Graph reuse and significantly reduce padding overhead.
@@
(b) Deadline-free mode: For offline tasks like dataset distillation lei2023comprehensive, each request does not have a preset deadline, and the policy reduces to token-max under the same feasibility constraints. The scheduler forms large, shape-similar short-prefill batches to fill the nearest captured CUDA Graph bucket (admit when $\mathrm{tok}(B)\geq M_{s}$), while long-prefill dispatches a single request with large fixed-size chunks $C_{l}$ to sustain high arithmetic intensity and maximize throughput.
Temporal disaggregation mode for single instance. PLA-Serve adopts a temporal disaggregation mode, where each GPU instance is dedicated exclusively to either short- or long-prefill execution. Two global queues $Q_{s}$ and $Q_{l}$ are maintained for short and long requests, respectively, and each instance pulls tasks only from its own queue. Scheduling decisions within each instance follow the policies described in the previous section: SLA-first (near-deadline priority) when deadlines are active, and token-max (CUDA Graph aggregation) when no deadline is preset. This exclusive-per-class execution avoids long-short interference and ensures stable prefill latency under both modes.
Spatial disaggregation mode for multiple instances. In the multi-instance setting, PLA-Serve employs a controller to dynamically balance short- and long-prefill workloads across $N$ GPU instances (see Algorithm 2). Two independent instance pools are maintained: $n_{s}$ short-prefill instances and $n_{l}=N-n_{s}$ long-prefill instances. At each control interval, the controller monitors the queue backlog, SLA deviation, and GPU utilization of each instance to estimate its load pressure. It then compares the overall pressures of the two pools and, after a cool-down period, migrates at most one instance between them when the imbalance exceeds a threshold. This simple feedback control stabilizes P99 latency, prevents oscillation, and keeps GPU utilization high with negligible overhead.

### v2

@@ Abstract
LAPS identifies and disaggregates requests with different prompt lengths in LLM serving to reduce TTFT latency. While recent systems have decoupled the prefill and decode stages to improve throughput, they still rely on unified scheduling policies that fail to adapt to heterogeneous workload characteristics. We observe that prompt-length variations lead to distinct performance bottlenecks, motivating an adaptive scheduling strategy. LAPS disaggregates multi-turn long-prefill requests from short-prefill ones and introduces a length-aware smart batching mechanism for short-prefill workloads. It adopts a dual-queue design that supports temporal disaggregation on a single prefill instance or spatial disaggregation across multiple instances. For short-prefill batches, a batch waiting window and CUDA Graph-based clustering mitigate interference from heterogeneous computation, reducing batching delay and lowering average latency. In real multi-turn workloads, LAPS reduces prefill latency by over 30% compared to vanilla SGLang under prefill–decode disaggregation, and further decreases SLO violations by 28% in multi-instance deployments with vanilla data-parallel configuration. Compared to the SGLang router with load balancing, it further lowers SLO violations by 12% in multi-GPU settings. Under high concurrency and mixed-request scenarios, LAPS improves request throughput by 35% serving Qwen2.5-32B model for prefill instance, demonstrating its effectiveness in optimizing heterogeneous LLM serving workloads.
@@ 1 Introduction
Modern LLM serving stacks (e.g., vLLM Kwon et al. (2023), SGLang Zheng et al. (2024b)) combine prefill-decoding (PD) disaggregation Zhong et al. (2024) with continuous batching to meet low-latency, high-concurrency service-level objectives (SLOs). The Prefill stage (first-token computation) is largely compute-bound, while the decoding stage (auto-regressive generation) is memory-bound. PD disaggregation decouples these two phases across separate instances to avoid cross-phase contention. On the prefill side, concurrent requests are batched to raise GPU utilization. However, we show that this separation and batching are insufficient: even with PD, interference still persists within the prefill stage when long, compute-bound prefills are mixed with short, memory-bound prefills/re-prefills.
Figure 1: P90 TTFT of long-prefill requests under varying concurrency levels for long and short requests. The long-prefill requests have more than 1K tokens, while the short ones have fewer than 64 tokens. We concurrently run them on a single H200 GPU and serve by Qwen2.5-32B Qwen et al. (2025). The dashed lines indicate the latency when only long-prefill requests are served.
@@
Re‑prefill denotes the repeated prefill in multi‑turn sessions where the model extends an existing context by combining new tokens with cached KV states. It is common in chatbots Dam et al. (2024), tool-using agents wölflein2025llmagentsmakingagent, RAG, speculative decoding Leviathan et al. (2023), and token routing She et al. (2025), and is typically memory-bound (dominated by KV-cache reads/writes rather than large GEMMs). Figure 2 illustrates the token length distribution of a real-world trace, LMsys-Chat-1M Zheng et al. (2024a), which is collected from real human-AI conversations. We observe that most prompts are short (<256 tokens), while long-context requests (>1K tokens) are relatively rare. This indicates that production workloads are dominated by short prefills/re-prefills, which are memory-bound. Mixing them with long, compute-bound prefills in unified batches will also induce compute-memory interference: short prefills/re-prefills wait behind long GEMMs and time-to-first-token (TTFT) spikes, and long prefills lose effective FLOPs due to heavy KV traffic from short jobs. Figure 1 confirms this issue: mixing long and short prefills does significantly increase long-prefill latency, and this contention becomes more severe as concurrency rises.
@@
Although prior systems have recognized the resource contention between compute-bound and memory-bound workloads, they focus on coordinating prefill and decode: (1) decode‑priority schedulersKwon et al. (2023) prioritize the decode phase to minimize per-token latency; (2) prefill-first schedulers prioritize the prefill phase and use continuous batching Yu et al. (2022) to improve throughput; (3) stall-free chunked prefill Agrawal et al. (2023) splits long prefill into chunks and interleaves them with decode, so long prefills don’t stall ongoing decodes; and (4) PD disaggregation Zhong et al. (2024); Jin et al. (2024); Hu et al. (2024a); Strati et al. (2024), which dominates in modern serving systems, runs prefill and decode on separate instances. These advances alleviate cross‑phase contention but implicitly assume all prefills are compute‑bound long sequences, overlooking that short (e.g., <256 tokens), memory‑bound prefills/re‑prefills could dominate real‑world multi-turn serving workloads.
@@
To address these challenges, we propose LAPS, a length-aware LLM serving system that explicitly disaggregates and optimizes long-prefill and short-prefill workloads within the prefill stage. LAPS maintains two separate prefill pools at runtime and performs batch disaggregation to isolate long and short prefill requests, completely eliminating their mutual interference. For short-prefill requests, LAPS introduces a dynamic waiting window in the scheduler and bucketizes requests by input length, making them more uniform and enabling larger batch sizes. This reduces batch-launch overhead and, combined with CUDA Graph-based execution, further accelerates processing and improves throughput. For SLO-serving scenarios, we design an SLO-aware scheduler that balances the trade-off between the waiting window and throughput.
@@
In multi-instance deployments, the scheduler dynamically adjusts each instance’s workload type based on real-time load, achieving load balancing across spatially disaggregated prefill instances. This adaptive strategy resembles resource allocation problems in deep learning systems, like in Qiao et al. (2021).
@@
In addition, LLM serving systems typically support three modes: mix and PD temporal/spatial disaggregation. Beyond these, our LAPS introduces a fourth mode:
@@
It is worth noting that our LAPS remains compatible with existing PD disaggregation architectures. Overall, our main contributions are summarized as follows:
@@ 2.1 Compute-Memory Boundary for (Re-)Prefills
Empirical profiling across advanced hardware generations (A100, H100, and H200) and LLMs ranging from 7B to 32B parameters shows that this transition typically occurs between 150 and 512 tokens Yuan et al. (2024); Zhong et al. (2024), depending on model architecture, kernel implementation, and batch configuration. Prefills shorter than this range tend to be memory-bound and limited by KV-cache I/O, while longer ones are dominated by GEMM throughput.
@@ 2.2 Exploring Interference Between Long-Short Prefills/Re-prefills
We model the interference between long-prefill and short-prefill (or re-prefill) requests using a standard M/G/1 (Markovian/General/1-server queue) first-come, first-served (FCFS) queuing model Meini (1998). Unlike prior PD scheduling analyses that focused on prefill-decode contention, we examine the intra-prefill interference caused by mixing compute-bound and memory-bound jobs within continuous batching.
@@
Using the Pollaczek-Khinchine (P-K) result Neuts (1986) for M/G/1 queues, the mean waiting time is
@@ 2.3 Uncovering the Sources of Long-Short Prefill Mixing
General-purpose LLM services must handle a wide spectrum of tasks. As shown in Table 1, daily chat and creative ideation are typical short-prompt tasks, while speculative decoding and token routing produce high-frequency, short re-prefills Chatterji et al. (2025). In contrast, long-document QA and autonomous agent workflows correspond to long-context prefills. In practice, these streams interleave over time, leading to long-short mixing within prefill workloads.
Most existing systems schedule requests in an FCFS fashion, packing them into unified batches. Many deploy multi-queue variants: continuous or rolling batching (e.g., vLLM Kwon et al. (2023), TGI HuggingFace (2024)) treats prefill and decode as distinct phases, applying FCFS-style admission under token/KV limits and optional priorities or aging policies. With chunked prefill (e.g., Sarathi-Serve Agrawal et al. (2023)), vLLM prioritizes decode and may co-batch prefills with decode. In-flight batching (e.g., TensorRT-LLM Corporation (2023)) runs distinct context and generation engines, each maintaining its own ready queue and often prioritizing generation. Moreover, many LLM gateways expose service tiers that prioritize certain request classes while enforcing token-based budgets using fixed or sliding windows (e.g., OpenAI OpenAI (2024), Anthropic Anthropic (2024), Envoy Proxy (2024), Kong Inc. (2024b), APISIX Foundation (2024), and Cloudflare Inc. (2024a)).
@@
The most related line of work is length bucketing, which groups requests by predicted sequence length into size-homogeneous buckets to reduce padding and improve throughput (e.g., Multi-Bin Batching Guldogan et al. (2024), BucketServe Zheng et al. (2025)). However, these methods only optimize intra-batch length variance; they do not disaggregate prefills versus re-prefills nor address the compute-memory interference we identify.
3 Length-Aware-Prefill Serving (LAPS)
@@
We develop LAPS to mitigate the interference between long-prefill and short-prefill requests in multi-turn LLM serving. The interference stems from two major factors: (1) their heterogeneous computation characteristics, and (2) head-of-line blocking caused by unified batching. Section 3.1 introduces the strategies we design to optimize high-concurrency short-prefill workloads, while Section 3.2 shows our queue- and instance-level disaggregation mechanism that isolates these request types to reduce interference. LAPS is built upon the prefill instance in the PD disaggregation architecture and extends it with a finer-grained disaggregation design within the prefill stage.
@@ 3.1 Short Prefill Optimization
During auto-regressive serving, the majority of end-to-end latency typically arises from the decode stage. Consequently, most existing optimization efforts (e.g., PD disaggregation Zhong et al. (2024), CUDA Graph acceleration Harish and Narayanan (2007), and router-based load balancing across decoding instances Hu et al. (2024b); Jain et al. (2025); Stripelis et al. (2024); Jitkrittum et al. (2025)) have been designed for the decode phase. However, as the diversity of LLM workloads grows (e.g., agent decision-making, chain-of-thought reasoning, and multi-turn task planning), the prefill stage has become an increasingly significant bottleneck, yet its optimization potential remains largely overlooked. Despite this trend, optimization for short and multi-turn prefills remains unexplored.
@@
Graph capture and bucketization. From the characteristics of intensive re-prefill workloads, speculative decoding and token-level routing—although not conventional PD inference—generate numerous short re-prefill requests. Drawing inspiration from EAGLE-2’s speculative decoding optimization Li et al. (2024; 2025), LAPS pre-defines a grid of power-of-two prompt-length-batch-size buckets (e.g., $L\in\{8,16,32,64,128,256\}$ and $B\in\{1,2,4,8,16,32,64\}$) Gao et al. (2023). At system initialization, a CUDA Graph is captured for each bucket under the assumption of fixed operator topology and variable memory addresses. During inference, each short-prefill request is padded to the nearest bucket and grouped with others sharing the same $(L,B)$ configuration, thereby maximizing graph reuse with negligible memory overhead.
Graph-aware memory-based batching. To maximize the benefit of CUDA Graphs for short-prefill workloads, LAPS optimizes batching with two goals: (i) reducing Graph launch frequency and (ii) increasing the reuse rate of large-batch Graphs. These are achieved through modestly extended waiting windows and graph fusion. Under high concurrency, slightly delaying batch formation allows more short-prefill requests to accumulate, improving overall efficiency when the saved launch overhead outweighs the waiting cost. Figure 5 shows the latency-throughput trade-off under different window settings.
@@
While current serving systems typically adopt a memory-constrained batching policy (i.e., aggregating requests until total tokens reach the GPU memory limit), LAPS enhances this approach with graph awareness. During short-prefill batching, requests are grouped under the memory budget and aligned to the nearest captured Graph shape.
@@ 3.2 Long-Short Prefill Disaggregation
To fundamentally eliminate the interference between long- and short-prefill requests discussed above, we adopt a design philosophy inspired by PD disaggregation, which further disaggregates long-prefill (LP) and short-prefill (SP) requests. However, unlike PD disaggregation, where the prefill and decode stages exhibit strong temporal dependencies and KV-cache transfers, the two types of tasks are merely mutually exclusive in our LP/SP disaggregation, resulting in fewer constraints in the scheduling objectives. Consequently, practical PD instance scheduling must account for the capacity coordination between the prefill and decode cluster, as well as the effective interconnect bandwidth when designing resource allocation strategies, but our design provides a larger design space for scheduling strategies that can adapt to the physical compute resources and workload characteristics. To this end, LAPS implements two complementary schedulers: a temporal disaggregation scheduler for single-instance prefill execution, and a spatial disaggregation scheduler for multi-instance prefill coordination. Figure 4 presents the system overview.
Disaggregating prefill execution eliminates direct interference between compute-bound long-prefill and memory-bound short-prefill tasks; however, static separation alone cannot accommodate dynamic workload variations. In real-world deployments, the ratio of long to short requests fluctuates over time, and requests within each category exhibit heterogeneous lengths and deadlines. To address this, LAPS introduces a hierarchical scheduling layer: a temporal disaggregation scheduler is employed within each single prefill instance to manage intra-instance prioritization, while a spatial disaggregation scheduler operates across multiple prefill instances to coordinate inter-instance resource allocation.
It is worth noting that the disaggregation design further amplifies the benefits of CUDA Graphs for short-prefill workloads. In mixed long/short prefill instances, a unified queue containing both request types limits the ability of short-prefill requests to form large, graph-aligned batches. In contrast, by maintaining two independent queues under the disaggregated design, LAPS can determine at request arrival whether CUDA Graph execution should be applied, thereby reducing batching delay and minimizing shape heterogeneity. In mixed queues, the large length disparity between long and short requests leads to excessive padding, lowering the Graph shape hit rate and GPU memory efficiency. After LP/SP separation, requests within each instance exhibit a more concentrated length distribution, improving both Graph reuse and throughput. As a result, the system can achieve higher CUDA Graph reuse and significantly reduce padding overhead.
@@
(b) Deadline-free mode: For offline tasks like dataset distillation Lei and Tao (2023), each request does not have a preset deadline, and the policy reduces to token-max under the same feasibility constraints. The scheduler forms large, shape-similar short-prefill batches to fill the nearest captured CUDA Graph bucket (admit when $\mathrm{tok}(B)\geq M_{s}$), while long-prefill dispatches a single request with large fixed-size chunks $C_{l}$ to sustain high arithmetic intensity and maximize throughput.
Temporal disaggregation mode for single instance. LAPS adopts a temporal disaggregation mode, where each GPU instance is dedicated exclusively to either short- or long-prefill execution. Two global queues $Q_{s}$ and $Q_{l}$ are maintained for short and long requests, respectively, and each instance pulls tasks only from its own queue. Scheduling decisions within each instance follow the policies described in the previous section: SLA-first (near-deadline priority) when deadlines are active, and token-max (CUDA Graph aggregation) when no deadline is preset. This exclusive-per-class execution avoids long-short interference and ensures stable prefill latency under both modes.
Spatial disaggregation mode for multiple instances. In the multi-instance setting, LAPS employs a controller to dynamically balance short- and long-prefill workloads across $N$ GPU instances (see Algorithm 2). Two independent instance pools are maintained: $n_{s}$ short-prefill instances and $n_{l}=N-n_{s}$ long-prefill instances. At each control interval, the controller monitors the queue backlog, SLA deviation, and GPU utilization of each instance to estimate its load pressure. It then compares the overall pressures of the two pools and, after a cool-down period, migrates at most one instance between them when the imbalance exceeds a threshold. This simple feedback control stabilizes P99 latency, prevents oscillation, and keeps GPU utilization high with negligible overhead.